what is the value of the expression 7x+2 if 5x-11=5
Answer by nyc_function(2741)

To find the value of 7x + 2, we first must solve 5x - 11 = 5 for x.
5x - 11 = 5

5x = 5 + 11
5x = 16

x = 16/5
We now replace x in the expression 7x + 2 with (16/5) and do the math.
7(16/5) + 2 = 122/5

Done!

Did you understand?